# RG evolution and effect of intermediate new-physics on $\Delta B=1$ four-fermion operators

### v1

@@ Abstract
Motivated by the stringent experimental bounds on proton lifetime and the need for precise low-energy predictions, there has been renewed interest in the renormalization group (RG) evolution of Wilson coefficients for baryon number violating (BNV) operators and their characteristic new-physics scales. In this work, we analyze the RG running of dimension-six four-fermion operators in the $\overline{\text{MS}}$ scheme that mediate nucleon decay channels such as $p\to e^{+}\pi^{0}$, while systematically accounting for the impact of baryon number conserving (BNC) new-physics that can enter the theory at an intermediate scale as higher-dimensional effective field theory operator. These BNC operators mix with BNV ones at 1-loop and alter the RG flow. The running is performed from the electroweak scale up to representative intermediate scales of $10^{4}~\text{GeV}$, $10^{6}~\text{GeV}$, and $10^{9}~\text{GeV}$, corresponding to possible thresholds for new BNC degrees of freedom. Comparing the RG evolved coefficients with current experimental bounds on nucleon decay lifetimes, we find that the inclusion of BNC-BNV mixing, dominated by top quark loops, can significantly lower the effective proton decay scale. A Python package [^4] Available at https://github.com/rp-winter/Nucleon-Decay-SMEFT is provided to facilitate the RG evolution of nucleon-decay Wilson coefficients, allowing for the inclusion of generic BNC effects.
@@ 1 Introduction
While baryon and lepton numbers appear as accidental global symmetries of the Standard Model (SM) at the classical level, they are explicitly broken by nonperturbative quantum effects [tHooft:1976rip], leaving $U(1)_{B-L}$ as the only conserved combination. Though there is no a priori reason to expect these symmetries to remain exact in BSM frameworks, many theoretical constructions, like Grand unified theories (GUTs) and related gauge extensions [Mohapatra:1980qe, Mohapatra:1980de, Mohapatra:1986dg, Mohapatra:1996pu, Babu:2008rq, Arnold:2012sd, Berezhiani:2015afa, Dev:2015uca, Allahverdi:2017edd, Fridell:2021gag, Gopalakrishna:2024qxk], naturally accommodate them. The relevant operators contributing to proton decay typically arise at dimension-six, and their coefficients are tightly constrained by the experimental limits on nucleon lifetimes (Table. 1), pushing the characteristic mass scale of these mediators close to $10^{15}$-$10^{16}$ GeV. The next generation of large-volume detectors, such as Hyper-Kamiokande [Hyper-Kamiokande:2018ofw], DUNE [2817610], and JUNO, are poised to extend the sensitivity of proton lifetimes to order $10^{35}$ years, offering unprecedented discovery prospects.
@@
While current experimental limits offer rough estimates of the energy scale associated with baryon number violation (BNV), these estimates are incomplete because of the significant gap between the low-energy scale of measurements and the GUT scale where new-physics matches onto the BNV operators. Ongoing efforts to incorporate next-to-leading order corrections, by including 2-loop renormalisation group (RG) evolution [Hisano:2013ege, Pokorski:2017ueo, Pokorski:2019ete] and dimension-7 operators [Beneito:2023xbk], are essential for translating experimental searches into better constraints on the landscape of viable BNV in BSM theories. Moreover, one must also account for loop-level mixing of possible intermediate scale baryon number conserving (BNC) new-physics. Consequently, if new BNC states exist in the intermediate scale range, that interact with light quarks, they are expected to generate loop-induced contributions to baryon number violating processes. This was studied in the context of $\Delta B=2$ processes in [ThomasArun:2025rgx], while, in here, we extend the analysis to $\Delta B=1$ processes. In contrast to previous approach, here we evolve the Wilson coefficient of proton decay operator with the beta function appended by the contribution from BNC dimension-six operators which mix with the BNV operators at 1-loop. A schematic of the analysis is given in Fig. 1. The loop diagrams (b) and (c) correspond to mixing in proton decay and mixing in the neutron decay processes respectively. Note that the contribution from light first two generation quarks, in the loop, is numerically suppressed. Although at first sight this contribution may appear to be suppressed by $\frac{m_{t}^{2}}{\Lambda^{4}}$, the insufficient experimental data on flavour violating BNC operator of the form $\overline{t}u\overline{t}u$’ and BNV operators like $ttde^{+}$’ allow for potential significant effects. While the analysis in this work focuses on the $p\to e^{+}\pi^{0}$ process, the framework can be effortlessly extended to other decay modes.
@@
To facilitate such further studies, we provide a publicly available Python package [RepoName] that performs the RG evolution. This tool is designed to flexible and easily accommodate a wide class of BNC and BNV Standard Model Effective Field Theory (SMEFT) new-physics operators at different choices of threshold energy scales, enabling users to generalize our analysis to other nucleon decays and to any BSM scenario they might wish to explore.
In Section. 2, we briefly introduce the SMEFT framework and the dimension-six baryon number conserving and baryon number violating operators. We also derive RG evolution of the proton decay operator, including the BNC-BNV mising at 1-loop. The running of the Wilson Coefficients are given in Section. 3, where we consider three scenarios corresponding to three different scales of BNC new-physics namely, $10^{4}$ GeV, $10^{6}$ GeV and $10^{9}$ GeV. The proton deacy Wilson coefficient is then evolved from $10^{3}$ GeV to these BNC energy sales. In Section. 4, we introduce the Python code that we used in this analysis and finally in Section. 5, we summarise our results. For readability, all the notations used in the paper are consolidated in Appendix B.
@@ 2 Theoretical Framework
In this section, we set up the theoretical framework required for our study. We begin with a brief overview of the SMEFT framework, with emphasis on dimension-six BNC and BNV operators, highlighting the subset of operators that mix at 1-loop. These BNC-BNV mixing is later used to study the modification to RG evolution of Wilson coefficients of operators that generate proton decay.
@@ 2.1 Standard Model Effective Field Theory
The Lagrangian at dimension-six operators takes the form,
@@
with $Q_{i,prst}$ denoting the complete, independent set of dimension-six operators and $c_{i,prst}$ their corresponding dimensionless Wilson coefficients. And prstprst, denotes the family indices of the four fermion fields. In our calculation we use dimentionful Wilson coefficients $C_{i,prst}$,
@@
These dimension-six operators have been classified into a complete and non-redundant basis. The lowest-dimensional BNV operators appear at dimension-six, where baryon number is violated by one unit ($\Delta B=1$). These operators play a central role in nucleon decay [Table. 1] and related processes.
@@
The BNC sector of SMEFT at dimension-six is remarkably rich. After using the equations of motion to eliminate redundant terms, there are 59 independent BNC operators. These operators are built exclusively from SM fields, preserve baryon and lepton number, and can be systematically grouped according to their field content:
@@
Although the number of operator structures is only 59, when one accounts for flavour indices, the parameter space expands enormously. For three generations, the complete dimension-six BNC Lagrangian contains 1350 CP-even and 1149 CP-odd parameters, yielding a total of 2499 real independent couplings. These parameters encode possible new-physics effects across a vast range of processes, including Higgs production and decay, electroweak precision observables, dipole transitions, and flavour physics.
@@
In the SMEFT framework, baryon number can be violated starting at dimension-six with four-fermion interactions involving three quarks and one lepton appear. These operators change baryon number by one unit ($\Delta B=1$), and lepton number by one unit ($\Delta L=1$), such that BLB-L remains conserved. They are the leading effective interactions responsible for nucleon decay in many ultraviolet completions of the SM, such as Grand Unified Theories (GUTs).
In the Warsaw basis, there are four independent $\Delta B=\Delta L=1$ operators at dimension-six:
@@
Altogether, the dimension-six SMEFT Lagrangian can therefore be schematically written as,
@@
The matching conditions for Wilson coefficients from the electro-weak scale (e.g. $10^{3}$ GeV) down to the hadronic scale are derived in the Ref. [Beneito:2023xbk]. Since in this work, our aim is to understand the involvement of dimension-six BNC operators lying at an intermediate scale between the electro-weak scale and BNV new-physics scale, we will adopt matching below electro-weak scale as in previous literature and do not repeat the discussion.
2.2 Loop Diagrams from BNV-BNC Mixing and Their Beta Functions
2.2.1 BNV-BNC Mixing
At the 1-loop level, the situation becomes richer. Insertions of the BNV operators can mix with SMEFT operators of the type $(\overline{q}q)(\overline{q}q)$, i.e. dimension-six four-fermion interactions, or with operators involving electroweak bosons and the Higgs sector. The “BNV block” in the Fig. 1 represents the proton decay operators, while the “BNC block” encapsulates the loop-induced contributions from the Standard Model fields and any possible intermediate new-physics degrees of freedom (for nucleon decay, these SMEFT contributions can only be BNC operators). Together, these blocks define the full EFT picture relevant for connecting high-scale baryon number violation with observable low-energy decay channels.
Among the complete set of 63 dimension-six SMEFT operators, only 22 operators enter into the loop-level BNV-BNC mixing for nucleon decay. The restriction arises because the fermionic structure of the BNV operators requires that the insertion from the BNC sector provide the correct spinor, colour, and gauge contractions to form a consistent 1-loop diagram. Operators that are purely bosonic, or that involve mismatched chirality structures, cannot close the loop with a BNV vertex.
@@ 2.2.1 BNV-BNC Mixing
Table 2: Baryon number conserving dimension-six SMEFT operators (22 in total) that contribute to 1-loop mixing with baryon number violating operators. The subscripts $p,r,s,t$ are flavour indices.
@@ 2.2.2 Beta Function Calculation
Where $g_{1}$, $g_{2}$, and $g_{3}$ are the gauge couplings of the Standard Model corresponding to the $U(1)_{Y}$ hypercharge, $SU(2)_{L}$ weak isospin, and $SU(3)_{c}$ colour interactions, respectively. The matrices $Y_{u}$, $Y_{d}$, and $Y_{e}$ denote the Yukawa couplings for up-type quarks, down-type quarks, and charged leptons, which enter through fermion-Higgs interactions and induce flavour-dependent operator mixing, and the subscripts $p,r,s,t,v,w$ are flavour indices. The coefficients $C^{\text{BNV}}_{i,prst}$ represent the Wilson coefficients of the dimension-six BNV operators, while $C^{\text{BNC}}_{j,prst}$ are those of the BNC SMEFT operators. Finally, $a_{ij}$ are the mixing coefficients introduced to parameterise the 1-loop contributions that connect the BNC and BNV sectors, extending the usual RG equations in which BNV operators only mix among themselves.
At tree level, the first-generation BNV operators directly mediate the decay $p\to\pi^{0}e^{+}$, which provides the leading experimental constraint on their Wilson coefficients [Beneito:2023xbk, Super-Kamiokande:2020wjk]. When loop effects are included, these same operators mix with BNC insertions through diagrams in which heavy quarks propagate in the loop. Among the possible choices, the top-top contractions give the dominant contribution to the anomalous dimensions, as illustrated in Fig. 2, which shows the Standard Model interaction together with the BNV-BNC mixing of the first-generation BNV operators. This implies that the most stringent bounds on the first-generation BNV coefficients receive loop-induced corrections controlled by the BNC sector, with the mixing coefficients $a_{ij}$ quantifying the strength of this connection. The explicit mapping between each proton decay BNV operator, the BNC and BNV operators with flavour violating structure that can mix into it, and the corresponding 1-loop coefficient is summarised in Table. 3. For completeness, the full calculation of a representative 1-loop diagram is presented in Appendix A.
Figure 2: Diagrammatic representation of BNV running (a) consists of tree (b), the SM interaction (c), and the BNV-BNC mixing (d) of first-generation BNV operators. The dominant loop contribution BNV-BNC mixing arises from top-top contractions, which feed into the anomalous dimensions of the BNV coefficients.
@@ 3 Running of the Beta Function
The running is performed for the scale $\Lambda/\sqrt{c}$, where $\Lambda$ represents the new-physics scale suppression and ‘$c$’ denotes the dimensionless Wilson coefficient of the effective operator defined in Eq. (3). This quantity serves as a convenient parameterisation of the effective interaction strength at different energy scales. Using the computed 1-loop beta functions from the previous section, the Wilson coefficients are evolved from $10^{3}$ GeV, electro-weak scale upward to the SMEFT scale, incorporating the full 1-loop contributions from both BNC and BNV sectors. Since the baryon number conserving (BNC) new-physics operator of the form $\overline{t}u\overline{t}u$ is not constrained to scales above $10^{4}~\text{GeV}$ by current Large Hadron Collider (LHC) data, we consider three representative scenarios with the BNC scale set at $10^{4}~\text{GeV}$, $10^{6}~\text{GeV}$, and $10^{9}~\text{GeV}$ for a comparative analysis.
3.1 Case 1: BNC operator scale at $10^{4}$ GeV

### v2

@@ Abstract
Motivated by the stringent experimental bounds on proton lifetime and the need for precise low-energy predictions, there has been renewed interest in the renormalization group (RG) evolution of Wilson coefficients for baryon number violating (BNV) operators and their characteristic new-physics scales. In this work, we analyze the RG running of dimension-6 four-fermion operators in the $\overline{\text{MS}}$ scheme that mediate nucleon decay channels such as $p\to e^{+}\pi^{0}$, while systematically accounting for the impact of baryon number conserving (BNC) new-physics that can enter the theory at an intermediate scale as higher-dimensional effective field theory operator. These BNC operators mix with BNV ones at 1-loop and alter the RG flow. The running is performed from the electroweak scale up to representative intermediate scales of $10^{4}~\text{GeV}$, $10^{6}~\text{GeV}$, and $10^{9}~\text{GeV}$, corresponding to possible thresholds for new BNC degrees of freedom. Comparing the RG evolved coefficients with current experimental bounds on nucleon decay lifetimes, we find that the inclusion of BNC-BNV mixing, dominated by top quark loops, can significantly lower the effective proton decay scale to $\sim 10^{7}$ GeV, thus mitigating the need of a large desert. A Python package [^4] Available at https://github.com/rp-winter/Nucleon-Decay-SMEFT is provided to facilitate the RG evolution of nucleon-decay Wilson coefficients, allowing for the inclusion of generic BNC effects.
@@ 1 Introduction
While baryon and lepton numbers appear as accidental global symmetries of the Standard Model (SM) at the classical level, they are explicitly broken by nonperturbative quantum effects [tHooft:1976rip], leaving $U(1)_{B-L}$ as the only conserved combination. Though there is no a priori reason to expect these symmetries to remain exact in BSM frameworks, many theoretical constructions, like Grand unified theories (GUTs) and related gauge extensions [Mohapatra:1980qe, Mohapatra:1980de, Mohapatra:1986dg, Mohapatra:1996pu, Babu:2008rq, Arnold:2012sd, Berezhiani:2015afa, Dev:2015uca, Allahverdi:2017edd, Fridell:2021gag, Gopalakrishna:2024qxk], naturally accommodate them. The relevant operators contributing to proton decay typically arise at dimension-6, and their coefficients are tightly constrained by the experimental limits on nucleon lifetimes (Table. 1), pushing the characteristic mass scale of these mediators close to $10^{15}$-$10^{16}$ GeV. The next generation of large-volume detectors, such as Hyper-Kamiokande [Hyper-Kamiokande:2018ofw], DUNE [2817610], and JUNO, are poised to extend the sensitivity of proton lifetimes to order $10^{35}$ years, offering unprecedented discovery prospects.
@@
While current experimental limits offer rough estimates of the energy scale associated with baryon number violation (BNV), these estimates are incomplete because of the significant gap between the low-energy scale of measurements and the GUT scale where new-physics matches onto the BNV operators. Ongoing efforts to incorporate next-to-leading order corrections, by including 2-loop renormalisation group (RG) evolution [Hisano:2013ege, Pokorski:2017ueo, Pokorski:2019ete, Naterop:2025lzc, Banik:2025wpi] and dimension-7 operators [Beneito:2023xbk], are essential for translating experimental searches into better constraints on the landscape of viable BNV in BSM theories. Moreover, one must also account for loop-level mixing of possible intermediate scale baryon number conserving (BNC) new-physics. Consequently, if new BNC states exist in the intermediate scale range, that interact with light quarks, they are expected to generate loop-induced contributions to baryon number violating processes. This was studied in the context of $\Delta B=2$ processes in [ThomasArun:2025rgx], while, in here, we extend the analysis to $\Delta B=1$ processes. In contrast to previous approach, here we evolve the Wilson coefficient of proton decay operator with the beta function appended by the contribution from BNC dimension-6 operators which mix with the BNV operators at 1-loop. A schematic of the analysis is given in Fig. 1. The loop diagrams (b) and (c) correspond to mixing in proton decay and mixing in the neutron decay processes respectively. Note that the contribution from light first two generation quarks, in the loop, is numerically suppressed. Although at first sight this contribution may appear to be suppressed by $\frac{m_{t}^{2}}{\Lambda^{4}}$, the insufficient experimental data on flavour violating BNC operator of the form $\overline{t}u\overline{t}u$’ and BNV operators like $ttde^{+}$’ allow for potential significant effects. While the analysis in this work focuses on the $p\to e^{+}\pi^{0}$ process, the framework can be effortlessly extended to other decay modes.
When all terms in the effective Lagrangian are taken into account, the BNV operators at dimension-8 are at the same order ($\mathcal{O}(\frac{1}{\Lambda^{4}})$) in expansion as the BNC-BNV mixing terms. Despite this similarity, their contributions to the RG evolution of the proton decay operator remains suppressed. To see this, recall that the dimension-8 BNV operators take the general forms [Murphy:2020rsh] $\psi^{4}H^{2}$, $\psi^{4}X$, $\psi^{4}D^{2}$, and $\psi^{4}HD$, where $H$, $X$, and $D$ denote the Higgs field, field-strength tensor, and covariant derivative, respectively. The contributions of these opeartors are either highly constrained, suppressed or appear at 2-loops compared to the BNC-BNV mixing effect. Nevertheless, they will induce other interesting processes such as $p\to e^{+}e^{+}e^{-}$, but this analysis lies beyond the scope of the present work.
@@
To facilitate further studies, we provide a publicly available Python package [RepoName] that performs the RG evolution. This tool is designed to flexible and easily accommodate a wide class of BNC and BNV Standard Model Effective Field Theory (SMEFT) new-physics operators at different choices of threshold energy scales, enabling users to generalize our analysis to other nucleon decays and to any BSM scenario they might wish to explore.
In Section. 2, we briefly introduce the SMEFT framework and the dimension-6 baryon number conserving and baryon number violating operators. We also derive RG evolution of the proton decay operator, including the BNC-BNV mising at 1-loop. The running of the Wilson Coefficients are given in Section. 3, where we consider three scenarios corresponding to three different scales of BNC new-physics namely, $10^{4}$ GeV, $10^{6}$ GeV and $10^{9}$ GeV. The proton deacy Wilson coefficient is then evolved from $10^{3}$ GeV to these BNC energy sales. In Section. 4, we introduce the Python code that we used in this analysis and finally in Section. 5, we summarise our results. For readability, all the notations used in the paper are consolidated in Appendix B.
@@ 2 Theoretical Framework
In this section, we set up the theoretical framework required for our study. We begin with a brief overview of the SMEFT framework, with emphasis on dimension-6 BNC and BNV operators, highlighting the subset of operators that mix at 1-loop. These BNC-BNV mixing is later used to study the modification to RG evolution of Wilson coefficients of operators that generate proton decay.
@@ 2.1 Standard Model Effective Field Theory
The Lagrangian at dimension-6 operators takes the form,
@@
with $Q_{i,prst}$ denoting the complete, independent set of dimension-6 operators and $c_{i,prst}$ their corresponding dimensionless Wilson coefficients. And prstprst, denotes the family indices of the four fermion fields. In our calculation we use dimentionful Wilson coefficients $C_{i,prst}$,
@@
These dimension-6 operators have been classified into a complete and non-redundant basis. The lowest-dimensional BNV operators appear at dimension-6, where baryon number is violated by one unit ($\Delta B=1$). These operators play a central role in nucleon decay [Table. 1] and related processes.
@@
The BNC sector of SMEFT at dimension-6 is remarkably rich. After using the equations of motion to eliminate redundant terms, there are 59 independent BNC operators. These operators are built exclusively from SM fields, preserve baryon and lepton number, and can be systematically grouped according to their field content:
@@
Although the number of operator structures is only 59, when one accounts for flavour indices, the parameter space expands enormously. For three generations, the complete dimension-6 BNC Lagrangian contains 1350 CP-even and 1149 CP-odd parameters, yielding a total of 2499 real independent couplings. These parameters encode possible new-physics effects across a vast range of processes, including Higgs production and decay, electroweak precision observables, dipole transitions, and flavour physics.
@@
In the SMEFT framework, baryon number can be violated starting at dimension-6 with four-fermion interactions involving three quarks and one lepton appear. These operators change baryon number by one unit ($\Delta B=1$), and lepton number by one unit ($\Delta L=1$), such that BLB-L remains conserved. They are the leading effective interactions responsible for nucleon decay in many ultraviolet completions of the SM, such as Grand Unified Theories (GUTs).
In the Warsaw basis, there are four independent $\Delta B=\Delta L=1$ operators at dimension-6:
@@
Altogether, the dimension-6 SMEFT Lagrangian can therefore be schematically written as,
@@
The matching conditions for Wilson coefficients from the electro-weak scale (e.g. $10^{3}$ GeV) down to the hadronic scale are derived in the Ref. [Beneito:2023xbk]. Since in this work, our aim is to understand the involvement of dimension-6 BNC operators lying at an intermediate scale between the electro-weak scale and BNV new-physics scale, we will adopt matching below electro-weak scale as in previous literature and do not repeat the discussion.
2.2 Loop Diagrams from BNC-BNV Mixing and Their Beta Functions
2.2.1 BNC-BNV Mixing
At the 1-loop level, the situation becomes richer. Insertions of the BNV operators can mix with SMEFT operators of the type $(\overline{q}q)(\overline{q}q)$, i.e. dimension-6 four-fermion interactions, or with operators involving electroweak bosons and the Higgs sector. The “BNV block” in the Fig. 1 represents the proton decay operators, while the “BNC block” encapsulates the loop-induced contributions from the Standard Model fields and any possible intermediate new-physics degrees of freedom (for nucleon decay, these SMEFT contributions can only be BNC operators). Together, these blocks define the full EFT picture relevant for connecting high-scale baryon number violation with observable low-energy decay channels.
Among the complete set of 63 dimension-6 SMEFT operators, only 22 operators enter into the loop-level BNC-BNV mixing for nucleon decay. The restriction arises because the fermionic structure of the BNV operators requires that the insertion from the BNC sector provide the correct spinor, colour, and gauge contractions to form a consistent 1-loop diagram. Operators that are purely bosonic, or that involve mismatched chirality structures, cannot close the loop with a BNV vertex.
@@ 2.2.1 BNC-BNV Mixing
Table 2: Baryon number conserving dimension-6 SMEFT operators (22 in total) that contribute to 1-loop mixing with baryon number violating operators. The subscripts $p,r,s,t$ are flavour indices.
@@ 2.2.2 Beta Function Calculation
Where $g_{1}$, $g_{2}$, and $g_{3}$ are the gauge couplings of the Standard Model corresponding to the $U(1)_{Y}$ hypercharge, $SU(2)_{L}$ weak isospin, and $SU(3)_{c}$ colour interactions, respectively. The matrices $Y_{u}$, $Y_{d}$, and $Y_{e}$ denote the Yukawa couplings for up-type quarks, down-type quarks, and charged leptons, which enter through fermion-Higgs interactions and induce flavour-dependent operator mixing, and the subscripts $p,r,s,t,v,w$ are flavour indices. The coefficients $C^{\text{BNV}}_{i,prst}$ represent the Wilson coefficients of the dimension-6 BNV operators, while $C^{\text{BNC}}_{j,prst}$ are those of the BNC SMEFT operators. Finally, $a_{ij}$ are the mixing coefficients introduced to parameterise the 1-loop contributions that connect the BNC and BNV sectors, extending the usual RG equations in which BNV operators only mix among themselves.
At tree level, the first-generation BNV operators directly mediate the decay $p\to\pi^{0}e^{+}$, which provides the leading experimental constraint on their Wilson coefficients [Beneito:2023xbk, Super-Kamiokande:2020wjk]. When loop effects are included, these same operators mix with BNC insertions through diagrams in which heavy quarks propagate in the loop. Among the possible choices, the top-top contractions give the dominant contribution to the anomalous dimensions, as illustrated in Fig. 2, which shows the Standard Model interaction together with the BNC-BNV mixing of the first-generation BNV operators. This implies that the most stringent bounds on the first-generation BNV coefficients receive loop-induced corrections controlled by the BNC sector, with the mixing coefficients $a_{ij}$ quantifying the strength of this connection. The explicit mapping between each proton decay BNV operator, the BNC and BNV operators with flavour violating structure that can mix into it, and the corresponding 1-loop coefficient is summarised in Table. 3. For completeness, the full calculation of a representative 1-loop diagram is presented in Appendix A.
Figure 2: Diagrammatic representation of BNV running (a) consists of tree (b), the SM interaction (c), and the BNC-BNV mixing (d) of first-generation BNV operators. The dominant loop contribution BNC-BNV mixing arises from top-top contractions, which feed into the anomalous dimensions of the BNV coefficients.
@@ 3 Running of the Beta Function
The running is performed for the scale $\Lambda/\sqrt{c}$, where $\Lambda$ represents the new-physics scale suppression and ‘$c$’ denotes the dimensionless Wilson coefficient of the effective operator defined in Eq. (3). This quantity serves as a convenient parameterisation of the effective interaction strength at different energy scales. Using the computed 1-loop beta functions from the previous section, the Wilson coefficients are evolved from $10^{3}$ GeV, electro-weak scale upward to the SMEFT scale, incorporating the full 1-loop contributions from both BNC and BNV sectors. Since the baryon number conserving (BNC) new-physics operator of the form $\overline{t}u\overline{t}u$ is not constrained to scales above $10^{4}~\text{GeV}$ by current Large Hadron Collider (LHC) data, we consider three representative scenarios with the BNC scale set at $10^{4}~\text{GeV}$, $10^{6}~\text{GeV}$, and $10^{9}~\text{GeV}$ for a comparative analysis. The initial values for both the BNC and BNV Wilson coefficients used in the analysis are summarised in Table. 4.